# Optimal Non-Adaptive Group Testing with One-Sided Error Guarantees

### v3

@@ IV Conclusion
It is shown in [16, Proof of Corollary 3.13] that if $T\leq\big(\frac{1}{\log 2}k\log_{2}k\big)(1-\varepsilon)$, and if there are $k=\tilde{n}^{\theta^{\prime}}$ defective items in a population of size $\tilde{n}$ for some $\theta^{\prime}$ sufficiently close to one such that $2(1-\theta^{\prime})<\xi<\varepsilon\theta^{\prime}$ for arbitrarily small $\xi>0$, the number of masked defective items $\widetilde{M}_{1}$ is stochastically dominated by $\mathrm{Bin}(\tilde{n}^{1-\xi},\frac{1}{3}\tilde{n}^{\varepsilon\theta^{\prime}-1})$. We then have
$\mathbb{P}\Big(\widetilde{M}_{1}\leq\frac{1}{4}\tilde{n}^{\varepsilon\theta^{\prime}-\xi}\Big)\leq\mathbb{P}\Big(\mathrm{Bin}\Big(\tilde{n}^{1-\xi},\frac{1}{3}\tilde{n}^{\varepsilon\theta^{\prime}-1}\Big)\leq\frac{1}{4}\tilde{n}^{\varepsilon\theta^{\prime}-\xi}\Big)=o(1)$ (95)
by standard binomial concentration (Appendix -A). Since rearranging $k=\tilde{n}^{\theta^{\prime}}$ gives $\tilde{n}=k^{\frac{1}{\theta}}$, we have
$\widetilde{M}_{1}\leq\frac{1}{4}\tilde{n}^{\varepsilon\theta^{\prime}-\xi}\iff\widetilde{M}_{1}\leq\frac{1}{4}k^{\varepsilon-\frac{\xi}{\theta^{\prime}}}.$ (96)
For context, we now outline how converse results are translated from high $\theta^{\prime}$ (near 1) to smaller $\theta$ in [16]. This is proved via the contrapositive statement that achievability for small $\theta$ implies the same for large $\theta^{\prime}$ (with the same number of tests, say $T=ck\log k$ for some $c>0$ when the two problems have the same $k$ but differing population sizes). To show this, the denser problem with parameter $\theta^{\prime}$ and $\tilde{n}$ items is solved by adding “dummy non-defectives” to increase the population size to $n\gg\tilde{n}$ (satisfying $n=\Theta(k^{\frac{1}{\theta}})$), without changing any of the test results, and the resultant sparser problem with parameter $\theta$ and $n$ items is then solved. Note that the full argument makes use of Lemma 12 and the set of satisfying sets $\mathcal{V}$} therein; the optimal MAP decoder samples uniformly from $\mathcal{V}$}, and this set is unchanged by the addition of dummy non-defectives.
We can use a similar idea for our problem setup. Namely, we can transfer our results from the high $\theta$ regime to the lower $\theta$ regime for $\theta\in\big(\frac{\log 2}{1+\log 2},\frac{1}{2}\big)$ by adding in these dummy non-defectives, and noting that adding in these non-defective items can only increase the number of masked non-defectives, and leaves the number of masked defective items unchanged (since non-defective items do not affect masking events of defectives). This means that $\widetilde{M}_{1}=M_{1}$, and (51) follows by combining this with (95)–(96).
-C2 Establishing Lemma 11
For Lemma 11, we will ultimately use the same “dummy non-defectives” idea as that of Lemma 10, but before doing so, we need to transfer a result from [30] from number of masked items to number of masked defectives.
We begin by using [30, Eqn. 3.12] and its preceding paragraph, which state that when using an i.i.d. prior with $\tilde{q}=\frac{k}{\tilde{n}}$ and $\tilde{n}$ items, where $k=\tilde{n}^{\theta^{\prime}}$ for $\theta^{\prime}=1-o(1)$, the number of masked defectives is stochastically dominated by $\mathrm{Bin}(\tilde{n}^{1-3\xi},e^{-\frac{(1+\xi)c_{\tilde{q}}}{\tilde{n}\tilde{q}}T})$ for some constant $c_{\tilde{q}}>0$ that depends on $\tilde{q}$. Moreover, when $\tilde{q}=o(1)$ (which is the case we consider), this constant is given explicitly by $c_{\tilde{q}}=(\log^{2}2)\big(1+o(1)\big)$. Hence, replacing $\tilde{n}\tilde{q}$ by $k$ in the denominator of the exponent and using standard binomial concentration bounds (Appendix -A), the number of masked items is at least
$\frac{1}{2}\tilde{n}^{1-3\xi}\exp\Big(-\frac{(1+\xi)\log^{2}2}{k}T\big(1+o(1)\big)\Big)$ (97)
with probability 1o(1)1-o(1). They then show the same holds for the combinatorial prior in [30, Section 3.2] by viewing the combinatorial prior as being generated in the following two-step procedure:
@@
This two-step procedure resembles the one used after Lemma 10, except we are now adding defective items from the first to the second step with high probability rather than removing them. The first step uses the i.i.d. prior, meaning we can apply (97); while $\tilde{q}^{\prime}$ here is slightly different from $\tilde{q}$ above, this only amounts to a change in the $1+o(1)$ term. The second step only consists of making more items defective, meaning that the number of masked items can only increase further, and (97) still holds. Moreover, under the high-probability event $k-2\sqrt{k}\log n\leq\lvert S\rvert^{\prime}\leq k$, the resultant distribution on $S$ follows the combinatorial prior.
As noted above, we seek to understand the number of masked defectives (rather than all masked items). Let $\widetilde{M}_{1}^{\prime}$ and $\widetilde{M}^{\prime}$ respectively denote the number of masked defective items and total masked items respectively between the first and second steps [^7] Specifically, we use this terminology to mean the number of defectives that would be masked if $S^{\prime}$ were the defective set. of the generation of $S$ described above. By the same reasoning as above, we have $\widetilde{M}_{1}^{\prime}\leq\widetilde{M}_{1}$. Moreover, we have $k-2\sqrt{k}\log\tilde{n}\leq\lvert S^{\prime}\rvert\leq k$ with probability 1o(1)1-o(1) due to the multiplicative Chernoff bound (Appendix -A), which implies that $\widetilde{M}_{1}$ equals the number of masked defectives under the combinatorial prior up to $o(1)$ probability events. Combining these observations, it is sufficient to show a high-probability lower bound on $\widetilde{M}_{1}^{\prime}$.
[…]
$\mathbb{E}[\widetilde{M}_{1}^{\prime}\mid\widetilde{M}^{\prime}=\tilde{m}^{\prime}$] $=\tilde{m}^{\prime}\tilde{q}^{\prime}$ (98)
$\geq\frac{1}{2}\tilde{q}^{\prime}\tilde{n}^{1-3\xi}\exp\Big(-\frac{(1+\xi)\log^{2}2}{k}T\big(1+o(1)\big)\Big)$ (99)
$=\frac{1}{2}\frac{k-\sqrt{k}\log\tilde{n}}{\tilde{n}}\tilde{n}^{1-3\xi}\exp\Big(-\frac{(1+\xi)\log^{2}2}{k}T\big(1+o(1)\big)\Big).$ (100)
We can then bound $\widetilde{M}_{1}^{\prime}$ below by
$\frac{1}{4}\cdot\frac{k}{\tilde{n}}\tilde{n}^{1-3\xi}\exp\Big(-\frac{(1+\xi)\log^{2}2}{k}T\big(1+o(1)\big)\Big)$ (101)
with probability 1o(1)1-o(1) due to binomial concentration bounds (Appendix -A). Since $\widetilde{M}_{1}^{\prime}\leq\widetilde{M}_{1}$, this then implies that
$\mathbb{P}\left(\widetilde{M}_{1}\leq\frac{1}{4}\cdot\frac{k}{\tilde{n}}\tilde{n}^{1-3\xi}\exp\Big(-\frac{(1+\xi)\log^{2}2}{k}T\big(1+o(1)\big)\Big)\right)=o(1).$ (102)
@@
Writing $\tilde{n}=k^{\frac{1}{\theta^{\prime}}}$ and applying some algebraic simplifications, (102) yields that with probability 1o(1)1-o(1), the number of masked defectives $\widetilde{M}_{1}$ is at least
@@
An outline of achievable approximate recovery rates using DD/COMP is given in [8, Section 5.1] for the case that $\eta^{-},\eta^{+}$ are fixed constants in $(0,1)$. However, we are unaware of any works giving full proofs, and moreover, the result that we seek in Theorem 1 also permits smaller values of $\eta^{-},\eta^{+}$, which is non-trivial to achieve. Thus, we provide a detailed proof for completeness, despite amounting to relatively straightforward extensions of existing work.
@@
[19, Lemma 1] Let ${\mathcal{M}}\subseteq[n]$ be an arbitrary set of items with $\lvert{\mathcal{M}}\rvert=M$, and let $W^{({\mathcal{M}})}$ be the number of tests in which at least one item of $\mathcal{M}$} is present. Fix constants $\alpha>0$ and $\delta>0$. Then, when drawing $LM$ tests with replacement for items in $\mathcal{M}$} to be included in, if $LM=\alpha T$ for some $\alpha=\Theta(1)$, then the following holds as $T\to\infty$:
@@
Recalling that $L=\frac{T\log 2}{k}$, we can apply this result with ${\mathcal{M}}=S$ and $\alpha=\log 2$, so that $W^{({\mathcal{M}})}$ corresponds to the number of positive tests. Hence, $W^{({\mathcal{M}})}=T_{1}$ concentrates around $\frac{T}{2}$ (i.e., $|T_{1}-\frac{T}{2}|\leq\delta T$ with probability 1o(1)1-o(1)). With this established, we can now proceed to bound the probability that a non-defective item only appears in positive tests. Fix some $i\in[n]\setminus S$; given $T_{1}$, each test that $i$ is drawn to be included in is positive with probability $\frac{T_{1}}{T}$. Thus we can bound the probability of $i$ only being included in positive tests as follows:
$\mathbb{P}(A_{i})$ $\coloneqq\mathbb{P}(i\ \mathrm{only\ in\ positive\ tests})$ (105)
@@
$=\sum_{t_{1}\,:\,t_{1}\leq(1+\delta)\frac{T}{2}}\mathbb{P}(T_{1}=t_{1})\left(\frac{t_{1}}{T}\right)^{L}+\sum_{t_{1}\,:\,t_{1}>(1+\delta)\frac{T}{2}}\mathbb{P}(T_{1}=t_{1})\left(\frac{t_{1}}{T}\right)^{L}$ (107)
$\leq\left(\frac{(1+\delta)T}{2T}\right)^{L}+2\exp\left(-\frac{\delta^{2}T}{\log 2}\right)$ (108)
$=\left(\frac{1+\delta}{2}\right)^{\frac{T\log 2}{k}}+2\exp\left(-\frac{\delta^{2}T}{\log 2}\right)$ (109)
$=\exp\left(-\frac{T\log 2}{k}\cdot\log\frac{2}{1+\delta}\right)+2\exp\left(-\frac{\delta^{2}T}{\log 2}\right),$ (110)
where (108) follows from Lemma 13 (and by bounding $\left(\frac{t_{1}}{T}\right)^{L}$ above by $\big(\frac{(1+\delta T)}{2T}\big)^{L}$ and $1$ in the first and second sums respectively, since it is an increasing function in $t_{1}$), and (109) follows from $L=\frac{T\log 2}{k}$. Taking $T=\frac{1}{\log 2\cdot\log\frac{2}{1+\delta}}k\log\frac{n}{k}=\big(\frac{1}{\log^{2}2}k\log\frac{n}{k}\big)(1+\varepsilon)$ for some arbitrarily small $\varepsilon>0$ (depending on arbitrarily small $\delta>0$), we obtain the following bound on $\mathbb{P}(A_{i})$:
$\mathbb{P}(A_{i})$ $\leq\exp\left(-(1+\varepsilon)\log\frac{n}{k}\right)+2\exp\left(-\frac{\delta^{2}T}{\log 2}\right)$ (111)
$=\left(\frac{k}{n}\right)^{1+\varepsilon}+2\exp\left(-\frac{\delta^{2}T}{\log 2}\right)$ (112)
$\leq 2\left(\frac{k}{n}\right)^{1+\varepsilon},$ (113)
where (113) holds for sufficiently large $n$. Therefore, denoting the number of non-defective items that only appear in positive tests as $H$, it follows that $\mathbb{E}H\leq 2n\left(\frac{k}{n}\right)^{1+\varepsilon}=2k\left(\frac{k}{n}\right)^{\varepsilon}$. By Markov’s inequality, we can then proceed to obtain a high probability bound on $H$:
$P_{\mathrm{e}}^{+}$ $=\mathbb{P}(H\geq\eta^{+}k)$ (114)
$\leq\frac{\mathbb{E}H}{\eta^{+}k}$ (115)
$\leq\frac{2k\left(\frac{k}{n}\right)^{\varepsilon}}{\eta^{+}k}$ (116)
$=k^{o(1)}\left(\frac{k}{n}\right)^{\varepsilon}$ (117)
@@
where (117) applies $\eta^{+}=k^{-o(1)}$, and (118) applies $k=\Theta(n^{\theta})$ with $\theta\in(0,1)$. Finally, we note that $T=\big(\frac{1}{\log^{2}2}k\log\frac{n}{k}\big)(1+\varepsilon)$ being sufficient for $P_{\mathrm{e}}^{+}\to 0$ implies that a rate of $\log 2$ is achievable for any $\theta\in(0,1)$. ∎
@@
Clearly, steps $1$ and $2$ make no mistakes, so any mistakes from the algorithm must be due to step $3$. In other words, any misclassification by the algorithm will be due to a defective item being marked as non-defective, implying that the estimate $\hat{S}_{\mathrm{DD}}$ returned by the algorithm satisfies $\hat{S}_{\mathrm{DD}}\subseteq S$. Thus, proving that the DD algorithm succeeds for SUBSET is equivalent to showing that the number of these false negative items is at most $\eta^{-}k$. To prove this for rates less than $\log 2$ (or equivalently for $T\geq\big(\frac{1}{\log 2}k\log_{2}\frac{n}{k}\big)(1+\varepsilon)$), we follow the steps of [19]. In particular we wish to show that
$\sum_{i\in S}\mathbb{P}(L_{i}=0)=o(\eta^{-}k),$ (119)
where $L_{i}$ is the number of tests in which $i$ appears and no other PD items appear. If we can show that this is the case, then this implies that the number of these false positive items is at most $\eta^{-}k$ with probability 1o(1)1-o(1) due to Markov’s inequality (since the above sum is the expected number of such items).
To do so, we first introduce some notation [^8] Due to notational clashes we use slightly different notation compared to [19], which denotes $\Phi$ as $\beta$. from [19] (again adopting the near-constant column weight design with $L=\frac{T\log 2}{k}$):

### v4

@@ IV Conclusion
The relevant analysis in [10] first bounds the number of masked defectives under the i.i.d. prior in a (relatively) dense setting, then translates it to the combinatorial prior in a dense setting, and finally to the combinatorial prior in a sparse setting. The details are given as follows.
Consider a setting with population size $\tilde{n}$ in which each item is defective independently with probability $\tilde{q}^{\prime}=\frac{k-\sqrt{k}\log\tilde{n}}{\tilde{n}}$, where $k=\tilde{n}^{\theta^{\prime}}$ for some $\theta^{\prime}$ sufficiently close to one such that $2(1-\theta^{\prime})<\xi<\varepsilon\theta^{\prime}$ for arbitrarily small $\xi>0$. Under this setting, it is shown in [10, Proof of Corollary 3.13] that the number of masked defectives $\widetilde{M}_{1}^{\prime}$ stochastically dominates $\mathrm{Bin}(\tilde{n}^{1-\xi},\frac{1}{3}\tilde{n}^{\varepsilon\theta^{\prime}-1})$. Accordingly, we have
$\mathbb{P}\Big(\widetilde{M}_{1}^{\prime}\leq\frac{1}{4}\tilde{n}^{\varepsilon\theta^{\prime}-\xi}\Big)$
$\leq\mathbb{P}\Big(\mathrm{Bin}\Big(\tilde{n}^{1-\xi},\frac{1}{3}\tilde{n}^{\varepsilon\theta^{\prime}-1}\Big)\leq\frac{1}{4}\tilde{n}^{\varepsilon\theta^{\prime}-\xi}\Big)=o(1),$ (114)
where the last step uses standard binomial concentration (Appendix -A). The analysis in [10] then relates this result to the number of masked defectives under the combinatorial prior (in a population of size $\tilde{n}$) via the following two-step generation of $S$:
@@
This two-step procedure resembles the one used after Lemma 10, except we are now adding defective items from the first to the second step rather than removing them. The first step uses the i.i.d. prior, meaning the high probability bound in (114) applies to the number of masked defectives in $S^{\prime}$. Since the second step involves making more items defective, the number of masked items following this step can only increase, meaning that $\widetilde{M}_{1}^{\prime}$ serves as a lower bound on the number of masked defectives under this distribution. Moreover, under the high-probability event $\{k-2\sqrt{k}\log n\leq\lvert S^{\prime}\rvert\leq k\}$, the resultant distribution on $S$ follows the combinatorial prior, meaning any bounds we derive under this distribution of $S$ also hold for the combinatorial prior up to $o(1)$ events.
Letting $\widetilde{M}_{1}$ denote the number of masked defectives under the above-mentioned combinatorial prior, the above discussion and (114) imply that $\widetilde{M}_{1}\geq\widetilde{M}_{1}^{\prime}\geq\frac{1}{4}n^{\varepsilon\theta^{\prime}-\xi}$ with probability 1o(1)1-o(1). Since rearranging $k=\tilde{n}^{\theta^{\prime}}$ gives $\tilde{n}=k^{\frac{1}{\theta^{\prime}}}$, we have
$\widetilde{M}_{1}\geq\frac{1}{4}\tilde{n}^{\varepsilon\theta^{\prime}-\xi}\iff\widetilde{M}_{1}\geq\frac{1}{4}k^{\varepsilon-\frac{\xi}{\theta^{\prime}}}.$ (115)
We now recap how converse results can be translated from denser settings to sparser settings, i.e., from higher $\theta^{\prime}$ (near 1) to smaller $\theta$. This is established in [10] via the contrapositive statement that achievability for small $\theta$ implies the same for large $\theta^{\prime}$ (with the same number of tests, say $T=ck\log k$ for some $c>0$ when the two problems have the same $k$ but differing population sizes). To show this, the idea is to start with a test design $\mathsf{X}$} for the sparser problem with parameters $n$ and $\theta$, and construct a test design $\mathsf{X}$$^{\prime}$ for the denser problem with parameters $\tilde{n}<n$ and $\theta^{\prime}>\theta$ as follows:
  1. 1.
    Choose a size-$\tilde{n}$ subset $V\subset[n]$ uniformly at random from the $n\choose\tilde{n}$} possibilities, and declare the remaining items $n]\setminus V$ to be (dummy) non-defectives.
  2. 2.
    Map the $\tilde{n}$ denser problem’s items to $V$ in an arbitrary fixed manner (e.g., in numerical order).
  3. 3.
    Perform the tests (and retrieve their results $\mathbf{Y}^{\prime}$) according to $\mathsf{X}$$^{\prime}$, defined to be the sub-matrix of the sparser problem’s test matrix $\mathsf{X}$} with the columns indexed by $V$.
We observe that since the items in $n]\setminus V$ are non-defective, the test results $\mathbf{Y}$ and $\mathbf{Y}^{\prime}$ generated by $\mathsf{X}$} and $\mathsf{X}$$^{\prime}$ respectively are the same (since the items removed to form $\mathsf{X}$$^{\prime}$ are designated as non-defective). Moreover, due to the symmetry of the construction, it is easy to see that if the denser problem has $k$ uniformly random defectives, then so does the sparser problem (accounting for the randomness in both the defective set and in $V$). Combining the findings from the preceding two sentences, we have that an achievability result for the sparser problem immediately implies the same for the denser problem. We refer the interested reader to [10] for the full details.
We can use a similar idea for our problem setup. Namely, we can transfer our results from the high $\theta$ regime to the lower $\theta$ regime for $\theta\in\big(\frac{\log 2}{1+\log 2},\frac{1}{2}\big)$ by adding in these dummy non-defectives, and noting that doing so can only increase the number of masked non-defectives, and leaves the number of masked defectives unchanged (since non-defective items do not affect masking events of defectives). This means that $\widetilde{M}_{1}=M_{1}$, and (69) follows by combining this with (114)–(115).
-C2 Establishing Lemma 11
Fix a test $t\in\{1,2,\dots,T\}$ containing $\gamma\geq\Gamma=\frac{n}{k}\log n$ items, and define the event ${\mathcal{E}}_{t}\coloneqq\{\lvert{\mathcal{I}}_{t}\cap S\rvert=0\}$ (i.e., the test outcome is negative), where we recall that ${\mathcal{I}}_{t}\subseteq[n]$ is the set of items included in test $t$. Without loss of generality, we assume that $\gamma=\Gamma$, since adding more items to the test can only decrease the probability of a negative test. Thus under the combinatorial prior, the probability of ${\mathcal{E}}_{t}$ occurring (with respect to the randomness in $S$) is bounded by
$\mathbb{P}({\mathcal{E}}_{t})$ $=\frac{\lvert\{s\in{\mathcal{S}}:\lvert{\mathcal{I}}_{t}\cap s\rvert=0\}\rvert}{{n\choose k}}$ (116)
$=\frac{{n-\Gamma\choose k}}{{n\choose k}}$ (117)
$=\frac{(n-\Gamma)(n-\Gamma-1)\cdots(n-\Gamma-k+1)}{n(n-1)\cdots(n-k+1)}$ (118)
$\leq\left(\frac{n-\Gamma}{n}\right)^{k}$ (119)
$\leq\exp\left(-\frac{k}{n}\Gamma\right)$ (120)
$=\frac{1}{n},$ (121)
where:
  • (116) uses $S\sim{\rm Unif}({\mathcal{S}})$;
  • (117) follows since any set $s$ such that $\lvert{\mathcal{I}}_{t}\cap s\rvert=0$ must be a subset of $n]\setminus{\mathcal{I}}_{t}$. Since $\lvert{\mathcal{I}}_{t}\rvert=\Gamma$, there are $n-\Gamma\choose k$} such sets;
  • (118) expands out the binomial coefficients and cancels terms;
  • (119) follows by upper bounding $\frac{n-\Gamma-i}{n-i}$ by $\frac{n-\Gamma}{n}$ for each $i=1,\dots,k-1$;
  • (120) uses $1-x\leq e^{-x}$ for $x\geq 0$, and applies this with $x=\frac{\Gamma}{n}$;
  • (121) follows by substituting the choice of $\Gamma$ and simplifying.
Taking a union bound over the at most $T=O(k\log n)=o(n)$ tests containing at least $\Gamma$ items yields the result.
Next, we show that if a test design $\mathsf{X}$} places at most $\Gamma$ items in each test, then the resultant number of items in at most $\log^{3}n$ tests is no(n)n-o(n). For each $i\in[n]$, define ${\mathcal{T}}_{i}\subseteq\{1,2,\dots,T\}$ to be the set of tests in which $i$ is included. A double counting argument then yields
$\sum_{i=1}^{n}\lvert{\mathcal{T}}_{i}\rvert=\sum_{t=1}^{T}\lvert{\mathcal{I}}_{t}\rvert\leq T\cdot\Gamma=O(n\log^{2}n),$ (122)
where the last inequality follows since $T=O(k\log n)$. Thus, the number of items that appear in more than $\log^{3}n$ tests is at most $O(\frac{n}{\log n})=o(n)$, which gives the desired result.
-C3 Establishing Lemma 12
For Lemma 12, we will ultimately use the same “dummy non-defectives” used in proving Lemma 10, but before doing so, we need to transfer a result from [5] from number of masked items to number of masked defectives.
We begin by using [5, Eq. 3.12] and its preceding paragraph, which state that when using an i.i.d. prior with parameter $\tilde{q}=\frac{k}{\tilde{n}}$ and $\tilde{n}$ items, where $k=\tilde{n}^{\theta^{\prime}}$ for $\theta^{\prime}=1-o(1)$, the number of masked items stochastically dominates $\mathrm{Bin}(\tilde{n}^{1-3\xi},e^{-\frac{(1+\xi)c_{\tilde{q}}}{\tilde{n}\tilde{q}}T})$ for some arbitrarily small $\xi>0$ and constant $c_{\tilde{q}}>0$ that depends on $\tilde{q}$. Moreover, when $\tilde{q}=o(1)$ (which is the case we consider), this constant is given explicitly by $c_{\tilde{q}}=(\log^{2}2)\big(1+o(1)\big)$. To obtain a similar result for masked defectives, it is useful to note how the binomial distribution arises in [5]:
  • Using the same ideas as [10], they identify a subset of $\tilde{n}^{1-3\xi}$ items that are masked with probability at least $e^{-\frac{(1+\xi)c_{\tilde{q}}}{\tilde{n}\tilde{q}}T}$, and whose masking events are independent;
  • The independence of masking events is ensured by a distance separation property, which can be stated as follows: For any two items $i$ and $i^{\prime}$ among the subset of size $\tilde{n}^{1-3\xi}$, the following two sets are disjoint: (1) items that share any test with $i$ (including $i$ itself); (2) items that share any test with $i^{\prime}$ (including $i^{\prime}$ itself).
To adapt the desired result from masked items to masked defectives, we simply observe that this separation property ensures that the property of being both masked and defective is still independent across these items. [^7] In fact, an equivalent property was implicitly used in [5, Eq. (3.14)] later in their analysis, and this was also used in [10] to establish the binomial distribution mentioned in the proof of Lemma 10 above. Combining this with the fact that any masked item is defective with probability matching the prior (i.e., $\tilde{q}$) [10, 5], it follows that the number of masked defectives stochastically dominates $\mathrm{Bin}(\tilde{n}^{1-3\xi},\tilde{q}\cdot e^{-\frac{(1+\xi)c_{\tilde{q}}}{\tilde{n}\tilde{q}}T})$. Observe that this matches the above-mentioned bound for masked items, but with the probability multiplied by $\tilde{q}$.
Using binomial concentration (Appendix -A) and $c_{\tilde{q}}=(\log^{2}2)\big(1+o(1)\big)$, it follows that the number of masked defectives (under an i.i.d. prior with probability $\tilde{q}=\frac{k}{\tilde{n}}$) is at least
$\frac{1}{2}\tilde{q}\cdot\tilde{n}^{1-3\xi}\exp\Big(-\frac{(1+\xi)\log^{2}2}{k}T\big(1+o(1)\big)\Big)$ (123)
with probability 1o(1)1-o(1), provided that (123) itself has $\omega(1)$ scaling (which will be true under our choice of $T$ below). Note that we have replaced $\tilde{n}\tilde{q}$ by $k$ in the exponent since the two are identical. Then, as noted in [5, Section 3.2], the exponent in (123) remains the same when the prior probability $\tilde{q}=\frac{k}{\tilde{n}}$ is replaced by $\tilde{q}^{\prime}=\frac{k-\sqrt{k}\log\tilde{n}}{\tilde{n}}$, since any changes are only in the $1+o(1)$ term. Letting $\widetilde{M}^{\prime}_{1}$ denote the number of masked defectives when each item is defective independently with probability $\tilde{q}^{\prime}$, we deduce that the following holds with probability 1o(1)1-o(1):
$\widetilde{M}_{1}^{\prime}$ $\geq\frac{1}{2}\tilde{q}^{\prime}\cdot\tilde{n}^{1-3\xi}\exp\Big(-\frac{(1+\xi)\log^{2}2}{k}T\big(1+o(1)\big)\Big)$ (124)
$\geq\frac{1}{4}\cdot\frac{k}{\tilde{n}}\tilde{n}^{1-3\xi}\exp\Big(-\frac{(1+\xi)\log^{2}2}{k}T\big(1+o(1)\big)\Big),$ (125)
where (125) uses the fact that $\frac{k-\sqrt{k}\log\tilde{n}}{\tilde{n}}\geq\frac{k}{2\tilde{n}}$ for sufficiently large $\tilde{n}$.
This bound can then be transferred to a bound on the number of masked defective items $\widetilde{M}_{1}$ under the combinatorial prior via the same argument involving the two-step generation of $S$ detailed in the proof of Lemma 10 above. We omit the details to avoid repetition, and simply write the resulting lower bound:
$\widetilde{M}_{1}\geq\frac{1}{4}\cdot\frac{k}{\tilde{n}}\tilde{n}^{1-3\xi}\exp\Big(-\frac{(1+\xi)\log^{2}2}{k}T\big(1+o(1)\big)\Big)$ (126)
@@
with probability 1o(1)1-o(1). Writing $\tilde{n}=k^{\frac{1}{\theta^{\prime}}}$ and applying some algebraic simplifications, (126) yields that with probability 1o(1)1-o(1), the number of masked defectives $\widetilde{M}_{1}$ is at least
@@
An outline of achievable approximate recovery rates using DD/COMP is given in [2, Section 5.1] for the case that $\eta^{-},\eta^{+}$ are fixed constants in $(0,1)$. However, we are unaware of any works giving full proofs, and moreover, the result that we seek in Theorem 1 also permits smaller values of $\eta^{-},\eta^{+}$, which is non-trivial to achieve. Thus, we provide a detailed proof for completeness, despite amounting to relatively straightforward extensions of existing work. As with Section III, we subsequently drop the superscripts from $\eta^{-},\eta^{+}$ since they will be clear from the context.
@@
[23, Lemma 1] Let ${\mathcal{M}}\subseteq[n]$ be an arbitrary set of items with $\lvert{\mathcal{M}}\rvert=M$, and let $W^{({\mathcal{M}})}$ be the number of tests in which at least one item of $\mathcal{M}$} is present. Fix constants $\alpha>0$ and $\delta>0$. Then, when drawing $LM$ tests with replacement for items in $\mathcal{M}$} to be included in, if $LM=\alpha T$ for some $\alpha=\Theta(1)$, then the following holds as $T\to\infty$:
@@
Recalling that $L=\frac{T\log 2}{k}$, we can apply this result with ${\mathcal{M}}=S$ and $\alpha=\log 2$, so that $W^{({\mathcal{M}})}$ corresponds to the number of positive tests. By doing so, we obtain that $W^{({\mathcal{M}})}=T_{1}$ concentrates around $\frac{T}{2}$ as follows:
$\mathbb{P}\left(\Big|T_{1}-\frac{T}{2}\Big|>\frac{\delta T}{2}\right)\leq 2\exp\left(-\frac{\delta^{2}T}{4\log 2}\right),$ (129)
where we rescaled $\delta$ to $\frac{\delta}{2}$ for later convenience. With this established, we can now proceed to bound the probability that a non-defective item only appears in positive tests. Fix some $i\in S^{\mathsf{c}}$; given $T_{1}$, each test that $i$ is drawn to be included in is positive with probability $\frac{T_{1}}{T}$. Thus, we can bound the probability of $i$ only being included in positive tests as follows:
$\mathbb{P}(A_{i})$ $\coloneqq\mathbb{P}(i\ \mathrm{only\ appears\ in\ positive\ tests})$ (130)
@@
$=\sum_{t_{1}=1}^{(1+\delta)\frac{T}{2}}\mathbb{P}(T_{1}=t_{1})\left(\frac{t_{1}}{T}\right)^{L}$
$\qquad+\sum_{t_{1}=(1+\delta)\frac{T}{2}+1}^{T}\mathbb{P}(T_{1}=t_{1})\left(\frac{t_{1}}{T}\right)^{L}$ (132)
$\leq\left(\frac{(1+\delta)T}{2T}\right)^{L}+2\exp\left(-\frac{\delta^{2}T}{4\log 2}\right)$ (133)
$=\left(\frac{1+\delta}{2}\right)^{\frac{T\log 2}{k}}+2\exp\left(-\frac{\delta^{2}T}{4\log 2}\right)$ (134)
$=\exp\left(-\frac{T\log 2}{k}\log\frac{2}{1+\delta}\right)+2\exp\left(-\frac{\delta^{2}T}{4\log 2}\right)$ (135)
where (133) follows from (129) (and by bounding $\left(\frac{t_{1}}{T}\right)^{L}$ above by $\big(\frac{(1+\delta)T}{2T}\big)^{L}$ and $1$ in the first and second sums respectively), and (134) follows from $L=\frac{T\log 2}{k}$. Taking $T=\frac{1+\delta}{\log 2\cdot\log\frac{2}{1+\delta}}k\log\frac{n}{k}=\big(\frac{1}{\log^{2}2}k\log\frac{n}{k}\big)(1+\varepsilon)$ for some arbitrarily small $\varepsilon>0$ (depending on arbitrarily small $\delta>0$), we obtain the following bound on $\mathbb{P}(A_{i})$:
$\mathbb{P}(A_{i})$ $\leq\exp\left(-(1+\delta)\log\frac{n}{k}\right)+2\exp\left(-\frac{\delta^{2}T}{4\log 2}\right)$ (136)
$=\left(\frac{k}{n}\right)^{1+\delta}+2\exp\left(-\frac{\delta^{2}T}{4\log 2}\right)$ (137)
$\leq 2\left(\frac{k}{n}\right)^{1+\delta},$ (138)
where (138) holds for sufficiently large $n$. Therefore, denoting the number of non-defective items that only appear in positive tests as $H$, it follows that $\mathbb{E}H\leq 2n\left(\frac{k}{n}\right)^{1+\delta}=2k\left(\frac{k}{n}\right)^{\delta}$. By Markov’s inequality, we can then proceed to obtain a high probability bound on $H$:
$P_{\mathrm{e}}^{+}$ $=\mathbb{P}(H\geq\eta k)$ (139)
$\leq\frac{\mathbb{E}H}{\eta k}$ (140)
$\leq\frac{2k\left(\frac{k}{n}\right)^{\delta}}{\eta k}$ (141)
$=k^{o(1)}\left(\frac{k}{n}\right)^{\delta}$ (142)
@@
where (142) applies $\eta=k^{-o(1)}$, and (143) applies $k=\Theta(n^{\theta})$ with $\theta\in(0,1)$. Finally, we note that $T=\big(\frac{1}{\log^{2}2}k\log\frac{n}{k}\big)(1+\varepsilon)$ being sufficient for $P_{\mathrm{e}}^{+}\to 0$ implies that a rate of $\log 2$ is achievable for any $\theta\in(0,1)$. ∎
@@
Clearly, steps $1$ and $2$ make no mistakes, so any mistakes from the algorithm must be due to step $3$. In other words, any misclassification by the algorithm will be due to a defective item being marked as non-defective, implying that the estimate $\hat{S}_{\mathrm{DD}}$ returned by the algorithm satisfies $\hat{S}_{\mathrm{DD}}\subseteq S$. Thus, proving that the DD algorithm succeeds for SUBSET is equivalent to showing that the number of these false negative items is at most $\eta k$. To prove this for rates less than $\log 2$ (or equivalently for $T\geq\big(\frac{1}{\log 2}k\log_{2}\frac{n}{k}\big)(1+\varepsilon)$), we follow the steps of [23]. In particular we wish to show that
$\sum_{i\in S}\mathbb{P}(L_{i}=0)=o(\eta k),$ (144)
where $L_{i}$ is the number of tests in which $i$ appears and no other PD items appear. If we can show that this is the case, then this implies that the number of these false positive items is at most $\eta k$ with probability 1o(1)1-o(1) due to Markov’s inequality (since the above sum is the expected number of such items).
To show (144), we first introduce some notation [^8] Due to notational clashes we use slightly different notation compared to [23], which denotes $\Phi$ as $\beta$. from [23] (again adopting the near-constant column weight design with $L=\frac{T\log 2}{k}$):